OpenStudy (anonymous):

solve for x. show the steps leading to the solution 13/2x-5=4/x

OpenStudy (mimi_x3):

13x = 8x - 20 5x = -20 x = -4

OpenStudy (anonymous):

is it \[13/(2x-5)=4/x\]

OpenStudy (anonymous):

ok \[13x=4(2x-5)\] \[13x=8x-20\] \[13x-8x=-20\] \[5x=-20\] \[x=-20/5\] \[x=-4\]